OpenStudy (anonymous):

Evaluate: 3(6 − 3) − 1 • 6

OpenStudy (ivettef365):

first solve the parenthesis

OpenStudy (jhannybean):

\[\large 3\color{green}{(6-3)}-1\cdot 6 \]\[\large 3\color{red}{(3)}-1 \cdot 6\]\[\large 9-6 =?\]

OpenStudy (anonymous):

yep^^